OpenStudy (unklerhaukus):

For the following DE, check that \(y_1\) is a solution of the corresponding homogeneous equation and then calculate the general solution by the reduction of order method. \[y^{\prime\prime} -4y^\prime +3y=3x-4;\qquad y_1 =e^x\]

OpenStudy (unklerhaukus):

\[y^{\prime\prime} -4y^\prime +3y=3x-4;\qquad y_1 =e^x\] \[\qquad\qquad\qquad\qquad\qquad\qquad y_1^\prime =e^x\]\[\qquad\qquad\qquad\qquad\qquad\qquad y_1^{\prime\prime} =e^x\]\[y_1^{\prime\prime} -4y_1^\prime +3y_1=0\]\[\left(e^x\right) -4\left(e^x\right)+3\left(e^x\right) =0\]\[\left(1-4+3\right)e^x=0\]%3x-4\]\[\therefore\qquad y_1= e^x\qquad\text{is a solution to the homogenous equation}\]\[y_2=uy_1=ue^x\]\[y_2^\prime=u^\prime e^x+ue^x=(u^\prime+u)e^x\]\[y_2^{\prime\prime}=(u^{\prime\prime}+u^\prime)e^x+(u^\prime+u)e^x\]\[y_2^{\prime\prime}=(u^{\prime\prime}+2u^\prime+u)e^x\]\[y_2^{\prime\prime} -4y_2^\prime +3y_2=3x-4\]\[\left((u^{\prime\prime}+2u^\prime+u)e^x\right) -4\left((u^\prime+u)e^x\right) +3\left(ue^x\right)=3x-4\]\[\left(u^{\prime\prime}+2u^\prime+u -4u^\prime-4u +3u\right)e^x=3x-4\]\[\left(u^{\prime\prime}-2u^\prime\right)e^x=3x-4\]\[u^{\prime\prime}-2u^\prime=\left(3x-4\right)e^{-x}\]\[\text{let } p=u^\prime\]\[p^\prime=u^{\prime\prime}\]\[p^\prime-2p=\left(3x-4\right)e^{-x}\]\[\mu(x)=e^{\int-2\text dx}=e^{-2x}\]\[\left(pe^{-2x}\right)^\prime=\left(3x-4\right)e^{-3x}\]\[pe^{-2x}=\int\left(3x-4\right)e^{-3x}\text dx\]\[pe^{-2x}=\int 3e^{-3x}x\text dx-4\int e^{-3x}\text dx\]\[pe^{-2x}=\left[-x{e^{-3x}}-\int -e^{-3x}\text dx\right]+\frac43e^{-3x}\]\[pe^{-2x}=-{xe^{-3x}}-\frac {e^{-3x}}3+\frac43e^{-3x}\]\[pe^{-2x}=-{xe^{-3x}}+e^{-3x}\]\[pe^{-2x}=(1-x)e^{-3x}\]\[p=(1-x)e^{-x}\]\[u^\prime=(1-x)e^{-x}\]\[u=\int(1-x)e^{-x}\cdot\text dx\]\[u=\int e^{-x}\cdot\text dx-\int xe^{-x}\cdot\text dx\]\[u=-e^x-\left[-xe^x-\int -e^{-x}\text dx\right]\]\[u=-e^{-x}+xe^x+e^{-x}\]\[u=xe^{-x}\]\[y_2=uy_1=xe^{-x}e^x=x\]\[y=ay_1+by_2\]

OpenStudy (unklerhaukus):

however the answer in the back of the book is \[y=ae^{x}+be^{3x}+x\] so i have made some mistakes/ left stuff out how do i incorporate the not homogenous part of the equation,

OpenStudy (anonymous):

hi UnkleRhaukus for Nonhomogeneous Differential Equations like \[y''+p(x) y'+q(x) y=r(x)\] the answer is \[y(x)=y _{c}+y _{p} \] Firstly, the 'yc' is the general solution to the homogeneous version of the equation. So you just write down the equation but with ZERO on the right-hand side, use the methods of the previous informations to find the solution, and that's your Complementary Function. The second part, 'yp' the particular integral, does involve the r(x) on the right-hand side. In fact, by looking at r(x), you have to guess what would be a likely form for the particular integral, and then try it in the equation.

OpenStudy (unklerhaukus):

hi mukushla what kind of guess to make ,

OpenStudy (unklerhaukus):

so\[y_c=ay_1+by_2\]\[y_c=ae^x+bx\]\[y=y_c+y_p\]\[y_p=?\]

OpenStudy (unklerhaukus):

e^3x?

OpenStudy (anonymous):

firstly for finding yc u must drop the (3x-4)

OpenStudy (unklerhaukus):

like this? \[y_c^{\prime\prime} -4y_c^\prime +3y_c=3x-4\] \[\left(D^2-4D+3\right)y_c=3x-4\]\[y_c=\frac{3x-4}{D^2-4D+3}\]

OpenStudy (unklerhaukus):

oh thats thats a different method to find \(y_p\) ;not right here.

OpenStudy (turingtest):

\[y''-4y'+3y=0\]is the homogeneous part

OpenStudy (anonymous):

yes...

OpenStudy (turingtest):

then you make a guess for the particular solution based on the RHS

OpenStudy (turingtest):

this method is undetermined coefficients guess that the particular solution is of the form\[y_c=Ax+B\]then plug that into the original DE and find the val;ues of A and B

OpenStudy (turingtest):

...or you could use another method: variation of parameters

OpenStudy (unklerhaukus):

which method will take the most lines of working

OpenStudy (turingtest):

oh I did neglect the fact that we may have to alter our guess for the particular depending on the solution to the complimentary, so you should find that first

OpenStudy (turingtest):

probably undetermined coefficients

OpenStudy (turingtest):

but first we should really get the homogeneous part

OpenStudy (unklerhaukus):

what have i done already?

OpenStudy (turingtest):

you must solve\[y''-4y'+3y=0\]you solved it with the 3x-4 part left in, which is wrong

OpenStudy (unklerhaukus):

oh yeah,

OpenStudy (turingtest):

you should be able to solve this pretty quick...

OpenStudy (unklerhaukus):

\[y_2^{\prime\prime} -4y_2^\prime +3y_2=0\]\[\left((u^{\prime\prime}+2u^\prime+u)e^x\right) -4\left((u^\prime+u)e^x\right) +3\left(ue^x\right)=0\]\[\left(u^{\prime\prime}+2u^\prime+u -4u^\prime-4u +3u\right)e^x=0\]\[\left(u^{\prime\prime}-2u^\prime\right)e^x=0\]

OpenStudy (unklerhaukus):

\[u^{\prime\prime}-2u^\prime=0\]

OpenStudy (unklerhaukus):

\[\text{let } p=u^\prime\]\[p^\prime=u^{\prime\prime}\]\[p^\prime-2p=0\]

OpenStudy (turingtest):

oh my bad, I didn't read that you had to use reduction of order...

OpenStudy (unklerhaukus):

\[\mu(x)=e^{\int-2\text dx}=e^{-2x}\]\[\left(pe^{-2x}\right)^\prime=0\]\[pe^{-2x}=\int0\cdot\text dx\]...

OpenStudy (unklerhaukus):

should i stop now,

OpenStudy (turingtest):

give me a minute to review the process and see if I can't still help...

OpenStudy (unklerhaukus):

ok,

OpenStudy (turingtest):

yeah, sorry I'm not sure how to do reduction of order with non-homogeneous DE's I could solve it, but not with that method.

OpenStudy (unklerhaukus):

ah, ok thanks for trying Turing test .

OpenStudy (turingtest):

good luck Unk

OpenStudy (anonymous):

what are you trying to do unkle

OpenStudy (anonymous):

i can help i just took a test on this

OpenStudy (anonymous):

\[y_2=y_1\int{}{}\frac{e^{-\int{}{}[p(x)dx}}{y_1^2}\]

OpenStudy (anonymous):

then use

OpenStudy (anonymous):

As you can see this is a pain in the butt so if you can use the reduction formula and the anti derivative comes out easy to do, i'd stick with it. If the antiderivative is something crazy. I'd do it the long way

OpenStudy (unklerhaukus):

where did you get that reduction formula @Outkast3r09

OpenStudy (anonymous):

@UnkleRhaukus ... let me grab my book

OpenStudy (unklerhaukus):

looks like an integrating factor in the numerator

OpenStudy (anonymous):

page 131 in the middle starts it... then i'll print the second page

OpenStudy (anonymous):

the reduction formula comes from seperation of variablees